OpenStudy (anonymous):

y varies jointly with p and q. When p = 4 and q = 5, y = 60. Find y when p = 5 and q = 8.

OpenStudy (anonymous):

y = k p q k = 60/20 = 3 y = 3 * 5 * 8 = 120

OpenStudy (anonymous):

@FutureMathProfessor You copied that from Yahoo answers

OpenStudy (anonymous):

Great